OpenStudy (anonymous):

plz help :( will give medal :)!<3 Several systems of equations are given below. System 1 y = 6x – 1.5 y = –6x + 1.5 System 1 x + 3y = –6 2x + 6y = 3 System 1 2x –y = 5 6x – 3y = 15 Which system of equations is consistent-independent? How many solutions will the system of equations have? Expain your answers. Which system of equations is consistent-dependent? How many solutions will the system of equations have? Expain your answers. Which system of equations is inconsistent-independent? How many solutions will the system of equations have? Expain your answers

OpenStudy (anonymous):

Each equation makes a rule. two simultaneous rules can a) make clear the only possible double-solution. b) have the same dependance to some variable.. c) contradict each other.

OpenStudy (anonymous):

are you familiar with elimination in systems of equations?

OpenStudy (anonymous):

a=intersecting lines / number of solution b=same line / many solution c=paraelle line / no solution

OpenStudy (anonymous):

yes ive learned it

OpenStudy (anonymous):

YES :) very true also a- consistent-independent b- consistent-dependent c- inconsistent-independent inconsistent: contradiction dependent: same line, same dependence

OpenStudy (anonymous):

y = 6x – 1.5 y = –6x + 1.5 wat do i do from here??

OpenStudy (anonymous):

graphing would be best, this is the first graph

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

thanks, that's a great graphing calculator :)

OpenStudy (anonymous):

haha thank ! : )np

OpenStudy (anonymous):

for the second graph, I re-arranged the variables like this: y = -2 -1/3 x y = 1/2 - 1/3 x

OpenStudy (anonymous):

x + 3y = –6 -2(x+3y=-6)? --> 2x + 6y = 3

OpenStudy (anonymous):

yes then can eliminate: -2x -6y = -6 2x + 6y = +3 -------------- 0 + 0 = -3 CONTRADICTION

OpenStudy (anonymous):

x + 3y = –6 -2(x+3y=-6)? --> -2x-6=12?

OpenStudy (anonymous):

true, it should be 12 -2x -6y = 12 2x + 6y = +3 -------------- 0 + 0 = +15 CONTRADICTION the lines of the equations in the second system are parallel and always contradict each other for any variable x

OpenStudy (anonymous):

i dont need to graph it?

OpenStudy (anonymous):

well you don't really have to, contradiction means they must be parallel can enter it into the graphic calculator nonetheless to double check

OpenStudy (anonymous):

mhhh i see ! : )

OpenStudy (anonymous):

2x –y = 5 <--(-3)--> (-6x+3y=-15) : D 6x – 3y = 15

OpenStudy (anonymous):

0 + 0 = 0, right? it's the exact same equation, once with factor 1 and once with factor 3

OpenStudy (anonymous):

it doesn't matter what x you choose, since they're compensated anyway

OpenStudy (anonymous):

yes o.o does that mean its B o.o

OpenStudy (anonymous):

yes :) it's B) - consistent (0 = 0) and dependent (on x: infinite solutions)

OpenStudy (anonymous):

wow that was kinda easy hehe ^_^ thanks so much for using your time with me !

OpenStudy (anonymous):

you're welcome